OpenStudy (anonymous):

a line has a rise of 6 and a slope of 1/2 find the run of the line

OpenStudy (johnweldon1993):

Well remember slope = rise / run So you would have \[\frac{ 6 }{ x } = \frac{ 1 }{ 2 }\] What value of 'x' makes this statement true?